Suppose x is a normal distribution random variable with mean 10 and standard deviation 1.5. Find a value of xo such that p(x>xo)=90%.

Group of answer choices

10+(-0.25)*1.5

10+1.28*1.5

10+(-1.28)*1.5

10+(-0.1)*1.5
Answer #1

Answer

Given that mean is 10 and standard deviation sd is 1.5

we have to find P(x>xo) = 90%

so, probability on lower side must be 10%

using percentile table for 10th percentile, we get z value = -1.28

So, required value of xo is given as

xo = mean + z*sd

= 10 + (-1.28)*1.5 = 8.08

therefore, option C is correct